Find the critical points and determine maxima or minima subject to the following constraint.
f(x,y,z) = 8x - 4z.
x^2+ 10y^2 + z^2 = 5.
1
Expert's answer
2020-06-14T17:40:52-0400

Apply the Lagrange multiplier method. Then "L(x,y,z,\lambda)=8x-4z+\lambda(x^2+10y^2+z^2-5)" then "L'_x=8+2\lambda x, L'_y=20\lambda y, L'_z=-4+2\lambda z" . Solving the system "L'_x=0,L'_y=0,L'_z=0"

and "x^2+10y^2+z^2=5" we get "x_1=-2,y_1=0,z_1=1, \lambda _1=2" and "x_2=2,y_2=0,z_2=-1,\lambda_2=-2" where "(-2,0,1),(2,0,-1)" is critical points.

Findiing "L''_{x^2}=2\lambda,L''_{y^2}=20\lambda,L''_{z^2}=2\lambda,L''_{xy}=L''_{xz}=L''_{zy}=0" we get

If "\lambda=2" then "(-2,0,1)" is point of minima constraint "f_{min}=-20"

If "\lambda=-2" then "(2,0,-1)" is point of maxima constraint "f_{max}=20"